# Degradation-Aware All-in-One Image Restoration via Latent Prior Encoding

### v2

@@ B.4 Downstream Vision Tasks (Object Detection)
Uformer Wang et al. (2022) 28.0 30.2 29.4 36.5 26.9 27.3
Restformer Zamir et al. (2022) 31.7 34.1 33.2 37.5 28.7 32.4
AIRNet Li et al. (2022) 30.2 32.7 31.8 35.8 26.4 30.3
PromptIR Vaishnav et al. (2023) 30.9 33.3 31.9 41.8 28.6 30.6
ADAIR Cui et al. (2024) 28.9 31.2 30.2 38.4 28.8 28.1
DFPIR Tian et al. (2025a) 27.9 30.0 29.3 26.6 25.5 29.8
@@
Table 10 presents a comprehensive performance comparison of various image restoration methods for object detection tasks. Our proposed DAIR method demonstrates superior performance across all evaluation metrics, achieving the highest Average Precision (AP) scores of 34.9% at IoU=0.50:0.95, 37.3% at IoU=0.50, and 36.4% at IoU=0.75, surpassing the second-best method (Restformer Zamir et al. (2022)) by a significant margin of 3.2 percentage points.DAIR also excels in detecting objects of varying sizes, achieving notably high AP scores on small (44.9%, outperforming PromptIR’s 41.8%), medium (30.8%), and large (35.1%) objects. The consistent improvements across all metrics highlight DAIR’s effectiveness in enhancing image quality for downstream object detection, with an average improvement of 3.2 percentage points over SOTA methods. These results demonstrate DAIR’s robust capability to address diverse degradation scenarios while preserving essential visual information for accurate object detection.
Appendix C Unseen Task Generalization
@@
A key motivation behind the proposed latent-prior DAIR framework is its ability to generalize to diverse, previously unseen degradations. To validate this capability, we extensively evaluate DAIR under cross-domain and out-of-distribution scenarios, including underwater image enhancement Peng & Cosman (2019), medical image perceptual enhancement Uhlen et al. (2010), medical image denoising Rezvantalab et al. (2018), real-world denoising Abdelhamed et al. (2018), and unseen real-world low-light enhancement Sharif et al. (2025). Notably, all evaluations are performed using a DAIR model trained only on common degradations, without any TS fine-tuning or external guidance. In addition to the visual results, we summarized the quantitative score with no-reference metrics: NIQE Mittal et al. (2013), MUSIQ Ke et al. (2021), and LPIPS Zhang et al. (2018b).
C.1 Real-World Denoising
@@
Table. 8 and Figure 9 illustrate the performance of the proposed method on real-world noisy images Abdelhamed et al. (2018). In well-exposed scenes, DAIR performs only denoising, preserving natural brightness and contrast. However, when inputs exhibit low-light characteristics (as commonly observed in SSID-like datasets Abdelhamed et al. (2018)) along with noise, the latent descriptor guides DAIR to jointly suppress noise and enhance brightness. Thus, DAIR adaptively performs denoising and enhancement without prompts or handcrafted rules.
Table 11: Image Quality Metrics for Real-world Noise Dataset. Best results in bold red.
Method Image Quality Metrics
NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
Input 5.57 61.31 0.81
Enhanced (Our) 4.24 69.37 0.80
Improvement +1.33 +8.06 +0.01
Figure 9: Real-world denoising results. Left: low-light noisy scenes where DAIR jointly performs denoising and LLIE automatically, without user input. Right: well-lit, noisy scenes where DAIR performs only denoising.
C.2 Underwater Image Enhancement
We further evaluate DAIR on underwater images, a domain unseen during training. As shown in Figure 10, the model adapts to varying underwater degradations, including haze-like scattering (left examples) and low exposure (right examples). Despite never being trained on underwater data, DAIR enhances visibility and color fidelity, demonstrating strong cross-domain generalization driven by its latent prior representation. Table 12 further details DAIR’s performance on underwater image enhancement.
Figure 10: Underwater image enhancement. DAIR generalizes to haze-like and low-exposed underwater degradations, improving contrast and color balance without domain-specific training.
Table 12: Image Quality Metrics for Underwater Image Enhancement. Best results in bold red.
Method Image Quality Metrics
NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
Input 3.83 53.35 0.70
Enhanced (Our) 3.78 54.05 0.70
Improvement +0.05 +0.70 +0.00
C.3 Unseen Low-Light Image Enhancement
We evaluate DAIR on real-world low-light images captured with unknown hardware and distributions (e.g., LSD-U Sharif et al. (2025)) unseen during training. Figure 11 and Table 13 demonstrate that the latent prior effectively encodes degradation properties, enabling robust out-of-distribution LLIE.
Figure 11: Unseen real-world low-light enhancement. DAIR can enhance illumination and structure in previously unseen low-light scenes without artifacts or over-enhancing the scenes.
Table 13: Image Quality Metrics for Real-world Lowlight Enhancement. Best results in bold red.
Method Image Quality Metrics
NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
Input 4.47 62.33 0.81
Enhanced (Our) 4.14 62.51 0.81
Improvement +0.33 +0.18 +0.00
C.4 Medical Image Perceptual Enhancement
We also study cross-domain perceptual enhancement on microscopic medical images Uhlen et al. (2010), which often suffer from poor contrast and weak structural saliency Sharif et al. (2022). Despite having no exposure to this modality during training, DAIR enhances perceptual quality by improving local contrast and clarity while preserving natural textures (Figure 12). These results demonstrate that the latent prior generalizes to structurally distinct domains. Table 14 presents the corresponding quantitative improvements.
Figure 12: Perceptual enhancement of microscopic medical images. DAIR improves contrast and structural visibility in unseen medical imagery without TS training.
Table 14: Image Quality Metrics for Perceptual Enhancement. Best results in bold red.
Method Image Quality Metrics
NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
Input 13.94 37.06 0.63
Enhanced (Our) 9.68 41.50 0.79
Improvement -4.26 +4.44 -0.16
C.5 Medical Image Denoising
Finally, we evaluate DAIR on medical image denoising using a dataset not seen during training, specifically dermatoscopic imagery Rezvantalab et al. (2018). As shown in Figure 13, DAIR effectively suppresses noise while preserving diagnostically relevant textures and edges, without requiring manual prompting. Table 15 illustrates the performance of DAIR on unseen medical image denoising.
Figure 13: Medical image denoising. DAIR reduces noise and preserves fine structural details in previously unseen medical data, operating without guidance or fine-tuning.
Table 15: Image Quality Metrics for Medical Noise Reduction. Best results in bold red.
Method Image Quality Metrics
NIQE $\downarrow$ MUSIQ $\uparrow$ LPIPS $\downarrow$
Input 16.99 21.93 0.82
Enhanced (Our) 12.64 25.17 0.70
Improvement +4.35 +3.24 +0.12
Appendix D Limitation and Future Scope
@@
Despite the significant improvement of our proposed DAIR method over existing approaches on both common and compound degradations, as well as its performance on unseen tasks, we observed certain limitations. Figure 14 illustrates some failure cases of our method on unseen tasks. Specifically, in extreme low-light conditions (under 5 lux), artifacts were observed in homogeneous spatial regions, such as cloud-free skies. Additionally, in underwater image enhancement, the method occasionally produced over-enhanced results in tricky scenes. These observations highlight areas that require further investigation, which we plan to address in future studies.

### v3

@@ D.1 Downstream Vision Tasks (OD)
Uformer 28.0 30.2 29.4 36.5 26.9 27.3
Restormer 31.7 34.1 33.2 37.5 28.7 32.4
AIRNet 30.2 32.7 31.8 35.8 26.4 30.3
PromptIR 30.9 33.3 31.9 41.8 28.6 30.6
ADAIR 28.9 31.2 30.2 38.4 28.8 28.1
DFPIR 27.9 30.0 29.3 26.6 25.5 29.8
@@
Table 20 presents a comprehensive performance comparison of various image restoration methods for OD tasks. Our proposed DAIR method demonstrates superior performance across all evaluation metrics, achieving the highest Average Precision (AP) scores of 34.9% at IoU=0.50:0.95, 37.3% at IoU=0.50, and 36.4% at IoU=0.75, surpassing the second-best method (Restformer (zamir2022restormer)) by a significant margin of 3.2 percentage points.DAIR also excels in detecting objects of varying sizes, achieving notably high AP scores on small (44.9%, outperforming PromptIR’s 41.8%), medium (30.8%), and large (35.1%) objects. The consistent improvements across all metrics highlight DAIR’s effectiveness in enhancing image quality for downstream OD, with an average improvement of 3.2 percentage points over SOTA methods. These results demonstrate DAIR’s robust capability to address diverse degradation scenarios while preserving essential visual information for accurate OD.
D.2 Unseen Task Generalization
@@
A key motivation behind the proposed latent-prior DAIR framework is its ability to generalize to diverse, previously unseen degradations. To validate this capability, we extensively evaluate DAIR under cross-domain and out-of-distribution scenarios, including underwater image enhancement (peng2017uieb), medical image perceptual enhancement (uhlen2010towards), medical image denoising (rezvantalab2018dermatologist), real-world denoising (abdelhamed2018ssid), and unseen real-world low-light enhancement (sharif2025illuminatingdarknesslearningenhance).
D.2.1 Comparison with existing Methods.
We also compared existing methods against DAIR on unseen real-world degradation. We benchmarked all models trained on common degradation without any task-specific fine-tuning. The quantitative scores are summarized using no-reference metrics: NIQE mittal2013niqe, MUSIQ ke2021musiq, and LPIPS zhang2018lpips. Table 21 presents a performance comparison on no-reference quality metrics. Our proposed method, DAIR, achieves the best results in most categories, demonstrating significant improvements, including a reduction of -0.78 in NIQE for unseen Real-Lowlight and an increase of +5.08 in MUSIQ. Furthermore, DAIR excels in Real-Noise and Real Perceptual Enhancement tasks, highlighting its effectiveness in enhancing image quality across a range of applications.
Table 21: Performance comparison across different image restoration tasks using no-reference quality metrics. Best results in bold red, second best underlined, and increment over best performing method highlighted in blue.
Method Real-Lowlight Real-Noise Real Underwater Real Perceptual Enhancement Medical Denoising Average
NIQE↓/MUSIQ↑/BRISQUE↓ NIQE↓/MUSIQ↑/BRISQUE↓ NIQE↓/MUSIQ↑/BRISQUE↓ NIQE↓/MUSIQ↑/BRISQUE↓ NIQE↓/MUSIQ↑/BRISQUE↓ NIQE↓/MUSIQ↑/BRISQUE↓
Input 5.33/47.30/27.17 4.44/62.56/5.55 3.83/52.35/8.80 13.94/37.06/56.26 17.06/21.81/65.10 8.92/44.42/32.58
PromptIR 5.31/42.58/39.07 5.77/61.11/27.38 3.98/54.00/17.45 7.59/42.62/38.14 25.60/18.72/78.09 9.65/43.81/40.03
DFPIR 5.50/45.81/21.83 4.88/59.72/16.59 4.06/52.16/8.97 9.55/43.13/15.52 16.04/24.27/64.71 8.01/45.02/25.52
ADAIR 5.53/46.71/35.77 4.38/63.69/16.30 3.63/55.50/13.29 11.08/40.27/33.54 9.65/24.34/43.36 6.85/46.10/28.45
DAIR (Ours) 4.75/51.79/21.97 4.24/69.37/16.94 3.61/55.32/15.50 5.42/45.94/34.91 8.22/28.53/24.55 5.25/50.19/25.97
Improvement -0.78/+5.08/+0.14 -0.14/+5.68/+0.35 -0.02/-0.18/+6.70 -2.17/+2.81/+19.39 -1.43/+4.19/-18.81 -1.60/+4.09/+0.45
D.2.2 Visual results
@@
Figure 11 illustrate the performance of the proposed method on real-world noisy images (abdelhamed2018ssid). In well-exposed scenes, DAIR performs only denoising, preserving natural brightness and contrast. However, when inputs exhibit low-light characteristics (as commonly observed in SSID-like datasets (abdelhamed2018ssid)) along with noise, the latent descriptor guides DAIR to jointly suppress noise and enhance brightness. Thus, DAIR adaptively performs denoising and enhancement without prompts or handcrafted rules.
Figure 11: Real-world unseen degradation restoration with DAIR. Our method can restore unseen out-of-distribution datasets without needing any fine tuning.
Similarly, our method can significantly improve the visual quality of unseen diverse real-world degradation, as shown in Fig. 12
Figure 12: Real-world unseen degradation restoration with DAIR. Our method can restore unseen out-of-distribution datasets without needing any fine tuning.
D.3 Real-world Compound Degradation
We evaluated our method on real-world compound degradation (i.e., LOL-Blur (zhou2022lednet). Table 22 presents a quantitative comparison of image restoration models on synthesized and real datasets. The results demonstrate that our proposed method, DAIR, outperforms existing models in all evaluated metrics. Notably, DAIR achieves the highest PSNR and SSIM values, as well as the lowest LPIPS and NIQE scores, indicating superior restoration quality compared to ADAIR. DAIR achives a PSNR of 24.10 and an SSIM of 0.8532 few-shot fine tunning. Fig. 13 further demonstrates the practicality of our method through a visual comparison.
Table 22: Quantitative comparison on synthesized and real datasets. Best results in bold red.
Model Training Synthesized Real
PSNR↑ SSIM↑ LPIPS↓ NIQE↓ MUSIQ↑ BRISQUE↓
Low (input) - 8.74 0.4255 0.4943 7.04 32.18 36.22
ADAIR Zero-shot 11.77 0.5528 0.4057 6.80 33.62 32.57
DAIR Zero-shot 13.35 0.6221 0.3393 5.14 46.39 28.58
ADAIR Fine-tune 22.83 0.7811 0.2406 6.80 33.62 32.57
DAIR Fine-tune 24.10 0.8532 0.1503 4.70 49.58 27.74
Figure 13: Real-world compound degradation restoration with DAIR.
Appendix E Limitation and Future Scope
@@
Despite the significant improvements of our DAIR method over existing approaches for both common and compound degradations, we identified some limitations. In real-world compound scenes, such as those in the LOL-Blur dataset(zhou2022lednet), our method can produce artifacts. Figure 14 shows failure cases on unseen tasks. Specifically, in extreme low-light conditions (under 5 lux), artifacts were observed in homogeneous spatial regions, such as cloud-free skies. Additionally, in underwater image enhancement, the method occasionally produced over-enhanced results in tricky scenes. These observations highlight areas that require further investigation, which we plan to address in future studies.